OpenStudy (anonymous):

Solve the equation on the interval 0≤θ<2π: tan(θ/2)=((√3)/3) Explain please!

hartnn (hartnn):

did you try ?

OpenStudy (anonymous):

I started with trying to separate tan(theta) from the rest of it, is that right?

hartnn (hartnn):

can you simplify this : ((√3)/3) ?

OpenStudy (anonymous):

I don't think so

hartnn (hartnn):

\(\sqrt a/a=\sqrt a/[(\sqrt a)(\sqrt a)]= 1/ \sqrt a\) did you get this ?

OpenStudy (anonymous):

So it would be 1/√3?

hartnn (hartnn):

yes, for what values of theta is tan value 1/sqrt 3 ??

OpenStudy (anonymous):

But wait, where did a or the denominator go

hartnn (hartnn):

'a' was just an example to show you that \(\sqrt3/3=1/\sqrt 3\) and which denominator ?

OpenStudy (anonymous):

The denominator of √3/3 o:

hartnn (hartnn):

denominator of √3/3 =3 but the whole √3/3 equals, 1/√3 so, tan (theta /2) = 1/ √3 for what values of angle is tan value 1/sqrt 3

OpenStudy (anonymous):

Oy okay. And I'm not sure o; do I have to solve out the different points on the unit circle to get it?

hartnn (hartnn):

tan = sin / cos so, see for which angle, this ratio comes out to be 1/sqrt 3

OpenStudy (anonymous):

π/3 and 4π/3?

hartnn (hartnn):

ok, range of theta is 0 to 2pi so, range of theta/2 (because thats our angle) will be 0 to pi/2 hence, we select only pi/3 theta /2 = pi/3 theta = .... ?

hartnn (hartnn):

range of theta/2 (because thats our angle) will be 0 to pi *****

OpenStudy (anonymous):

π/3? o:

hartnn (hartnn):

theta /2 = pi/3 theta = 2pi/3 o:

OpenStudy (anonymous):

And that's the answer? o;

hartnn (hartnn):

yes, thats it.

hartnn (hartnn):

any doubts ?

OpenStudy (anonymous):

My teacher put down that the answer was π/3, did I mess up somewhere? o:

hartnn (hartnn):

oh, yes :P the angles are not π/3 and 4π/3 for tan value of 1/sqrt 3 i didn't check/verify, assumed you are correct :P

hartnn (hartnn):

its pi/6 in the range 0 to pi so, theta /2 = pi/6 theta = 2pi/6 = pi/3

hartnn (hartnn):

sin pi/6 = 1/2 cos pi/6 = sqrt 3/ 2 tan pi/6 = sin pi/6 / cos pi/6 = 1/ sqrt 3

OpenStudy (anonymous):

Oh okay, I got it! Thank you very much (:

hartnn (hartnn):

welcome ^_^